TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x05x04x16x03

Cooldown Sector 78

Despite the long series of faults (network, cooling water, pumping group) experienced at the end of last week, the cryo team managed to bring down the average temperature of the magnets to 3 K during the week-end. The switch over to the new cryoplant took place on Monday but the connection of the new pumping group was delayed until Wednesday morning. Stabilization of temperatures and pressures as well as fine adjustments of the plants took place most of Wednesday until Thursday at 11.30 when the pumping group was set at 15 mbar. Unfortunately, at 11.30 an attempt to restart the UPS system on the surface caused a power cut which stopped all the PLCs. The cryo team is still recovering from that failure and we expect to be at nominal cryogenic condition in the arc tonight Friday May 4th.

Following a top management meeting which took place on Monday, a series of measures towards the clearing of the difficulties with the commissioning of the DFBs were taken: two different ways to avoid condensation on the chimneys are now in place: blowers on 6kA leads + heaters on the 120A leads on the DFBMC and dry-air bags around the chimneys of the DFBMA.

Although the powering front was very quiet during the week, significant progress was achieved with the preparation of the sequencer for the independently powered quadrupoles and the 120A correctors: we now have the sequencer tuned for the 60 A and 120 A correctors as well as the independently powered quadrupoles and D2.


Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-04 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login